OpenStudy (mtalhahassan2):

Adding and subtracting Rational Expression 2m divide by m-9 + 18 divide by 9-m

OpenStudy (anonymous):

\[18/(9-m)=-18/(m-9)\] \[2m/(m-9) -18/(m-9)=2(m-9)/(m-9)=2 \]